OpenStudy (anonymous):

Jina needs to memorize words on a vocabulary list for Korean class. She has memorized 24 of the words, which is two-thirds of the list. How many words are on the list?

OpenStudy (aaronandyson):

Let x be the total number of words on the list.

OpenStudy (aaronandyson):

She read 2/3 of the words so, \[\frac{ 2 }{ 3 } \times x = 24\]

OpenStudy (aaronandyson):

So now 2 times x = 24 times 3

OpenStudy (aaronandyson):

Solve for x now.

OpenStudy (aaronandyson):

\[2 \times x = 24 \times 3\]

OpenStudy (anonymous):

i got 72 ???

OpenStudy (aaronandyson):

2 times x = 72

OpenStudy (aaronandyson):

so now x = 72/2

OpenStudy (anonymous):

i got it by doing 24 * 3

OpenStudy (anonymous):

so 36 ?

OpenStudy (aaronandyson):

CORRECT!!!

OpenStudy (anonymous):

yay thank you